please help solve equation 8x + -2 = -9 + 7x. solution needs to be explained to a 6th grader that is having a difficult time getting it.
Answer by Alan3354(69443)

8x + -2 = -9 + 7x
----------------------------
Do the same thing to both sides of the equal so it remains equal.
8x + -2 = -9 + 7x
Subtract 7x
1x - 2 = -9
x - 2 = -9
Add 2
x = -7
We're done, have a beer.